OpenStudy (anonymous):

which of the following equations if equivalent to the given equation? 1/2x-6y=10 A.y=3y+5 B.x=12y+10 C.y=3x-5/3 D.y=1/12x-5/3

OpenStudy (smarty):

B

OpenStudy (anonymous):

D

OpenStudy (anonymous):

-6y=-1/2x+10 y=1/12x-5/3